OpenStudy (anonymous):

Two forces act on a hook. The magnitudes are F1= 300N and F2= 125N. The angle between the two forces is 48. Find the direction(with respect to F1) and the magnitude of the resultant.

OpenStudy (anonymous):

 R^2=P^2+Q^2+2PQ Cos x tany=Q Sinx/(P+Qcosx),where y is the angle between P&R F^2=300^2+125^2+2(300)(125)COS48 F^2=25^2[12^2+5^2+2*12*5Cos48] F^2=625[144+25+120Cos48]=625(169+120cos48) F=25 sqrt(169+120cos48) If y is the angle between F1 &F tan y=125 sin48/(300+125cos48) i have no calculator,you can solve cos48& sin48

OpenStudy (anonymous):

TakeF=R

OpenStudy (anonymous):

Thank You so much!